### v1

@@ 4 Related Works
Bucarey et al. (2022) proposed an algorithm for “strong” SSEs in the 2-player general-sum stochastic game. Assuming the uniqueness of the follower’s best response, the problem setting of Bucarey et al. (2022) is the same as ours. However, their algorithm requires strong assumptions to guarantee convergence to the SSEs, and there is little assurance of the performance of the obtained policy when it does not converge to the SSEs or when there is no SSE. This problem is because the fixed point of the operator used in Bucarey et al. (2022) is different from the SE value function. We discuss this in detail in Section 5.1.
@@
A case where the algorithm of Bucarey et al. (2022) converges to the SSE is when the follower is myopic, which maximizes its immediate reward. There are some existing studies in the SSGs under the myopic follower. For example, Zhao et al. (2023) derived the upper bound of the regret of the leader’s payoff in the cooperative tasks, where the reward function is shared between the leader and the follower, with information asymmetry under the myopic follower. Zhong et al. (2023) proposed the guaranteed algorithm in the SSG where the group of myopic followers forms a Nash equilibrium among them. Different from these works, we focus on the situation where the follower is not myopic.
Some studies concentrate on the cooperative setting. Kao et al. (2022) and Zhao et al. (2023) analyzed the cooperative tasks with information asymmetry, where the leader cannot observe the follower’s action Kao et al. (2022) and the reward functions are known only to the follower Zhao et al. (2023). Kononen (2004) and Zhang et al. (2020a) point out Pareto efficiency and uniqueness as advantages of Stackelberg equilibria to Nash equilibria in cooperative tasks.
@@
Zhang et al. (2020a) is similar to ours and Bucarey et al. (2022) in that the problem setting is the general-sum and the follower is not myopic. The differences are that the transition function is unknown (i.e., in the RL setting), the follower’s policy takes the leader’s action as input, and only deterministic stationary policies are considered the leader’s policies. The algorithm proposed by Zhang et al. (2020a) shares its principal foundation with that of Bucarey et al. (2022) (see Appendix L). This means that these methods use the same operator, implying that the limitation of the method of Bucarey et al. (2022) applies to that of Zhang et al. (2020a).
We proposed a novel algorithm for general-sum SSGs where the transition function and the reward functions are known and the follower is not myopic, which is the same as Bucarey et al. (2022). Different from the previous work, we proposed a novel approach with a convergence guarantee even when no SSE policy exists to deal with such a case.
5 Fixed Point Analysis of Operators
5.1 Operator defined by Bucarey et al. (2022)
Bucarey et al. (2022) addressed the same problem as ours under the uniqueness of $R_{B}^{*}$ [^1] Bucarey et al. (2022), which defines Strong SSEs to avoid the ambiguity of the best responses and utilizes it as the optimal solution of the problem. In this study, instead, we assume that there is no ambiguity in the best responses; hence, the SSEs are indeed strong SSEs. . They defined an operator whose fixed point coincides with the SE value function under certain conditions. Then, they proposed a method to construct a stationary equilibrium policy corresponding to the fixed point of the defined operator, which is called fixed point equilibrium (FPE).
However, they required strong assumptions that can potentially limit the applications of their approach (see for example Appendix M). First, they assumed that the follower is “myopic” or the reward functions are either cooperative or competitive to guarantee the convergence of their algorithm to the FPE Bucarey et al. (2022); Zhang et al. (2020a). Second, to guarantee the coincidence of the FPE and the SSE, the myopic follower was assumed Bucarey et al. (2022). Notably, the situation where the follower is “myopic” in Bucarey et al. (2022) is equivalent to the one where the follower’s discount rate $\gamma_{B}=0$ or the one where the transition function is not dependent on the follower’s actions Bucarey et al. (2022). That is, if the follower is not myopic, the algorithm may not converge and the fixed point of the algorithm may be different from the SSE.
Even if the convergence of the algorithm is not guaranteed, it is of great significance from a practical viewpoint to understand the characteristics of the FPE if exists. Particularly, it is interesting to examine how effective the FPE is if it does not coincide with the SSE. Thus, we analyzed the fixed point of the operator defined in Bucarey et al. (2022).
The operator $T:\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$ defined by Bucarey et al. (2022) for the strong SSE policies can be simplified as one for the SSE policies as: for $v:=(v_{A},v_{B})\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$, $i\in\{A,B\}$ and $s\in\mathcal{S}$,
$(Tv)_{i}(s):=r_{i}(s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))\\ +\gamma_{i}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))}\left[v_{i}(s^{\prime})\right],$ (2)
where $R_{A}(s,v)$ is
@@ 5 Fixed Point Analysis of Operators
$R_{A}(s,v):=\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},v_{B}))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},v_{B}))}\left[v_{A}(s^{\prime})\right],$
@@
and is assumed to exist for all $s$ and $v$, [^2] This assumption is necessary to derive a policy corresponding to a value function and Bucarey et al. (2022) implicitly assume it. However, it is not guaranteed because the inside of $\operatornamewithlimits{argmax}$ is not necessarily continuous with respect to $f_{s}$, even though $\Delta(\mathcal{A})$ is compact. $R_{B}$ always exists because $\mathcal{B}$ is finite, and
@@
$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)\\ +\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right]$
@@
is assumed to be a singleton for all $s$, $f_{s}$, and $v_{B}$. [^3] Bucarey et al. (2022) defines the “strong” version of $R_{B}$ similarly to the strong SSE. Thus, we assume that $R_{B}$ is a singleton to simplify the discussion in accordance with our goal to find the SSE policies. As the core of our analysis does not depend on the uniqueness of $R_{B}$, the result can be easily extended to the method of Bucarey et al. (2022).
@@
If there exists a fixed point $V:=(V_{A},V_{B})$ of $T$, the pair of the value functions of the FPE $(\bar{f},\bar{g})$ defined as follows coincides with $V$ Bucarey et al. (2022): for any $s\in\mathcal{S}$,
$\bar{f}(s)\in R_{A}(s,V),$ $\bar{g}(s)\in R_{B}(s,\bar{f}(s),V_{B}).$
$T$ is a contraction mapping for $v_{A}$ under certain conditions [^4] Bucarey et al. (2022). An example is the case when $r_{A}$ is an affine transformation of $r_{B}$. In this situation, the game is either cooperative or competitive.. Then, we can obtain the fixed point asymptotically by repeatedly applying such $T$ to an arbitrary initial function $v_{A}^{(0)}\in\mathcal{F}_{\mathcal{S}}$ Bucarey et al. (2022). In this case, a single update by $T$ can be interpreted as solving a normal-form Stackelberg game on each state $s\in\mathcal{S}$ by considering the current Q functions
$q_{i}^{(t)}(s,a,b):=r_{i}(s,a,b)+\gamma_{i}\mathbb{E}_{s^{\prime}}\left[v_{i}^{(t)}(s^{\prime})\right]\quad(i\in\{A,B\})$
as the gain tables on $s$ for the leader and the follower. This operation is also used in Nash-Q Learning Hu & Wellman (2003), which is guaranteed to converge to the Nash equilibrium in stochastic games. However, the convergence to the Stackelberg equilibrium by the same operation is not guaranteed even if the equilibrium exists.
@@
Let us consider the characteristics of the fixed point $(V_{A},V_{B})$ of the operator $T$ in general situations where the fixed point exists but it is different from the SE value functions. Theorem 5.1 reveals a general property of $(V_{A},V_{B})$. The proof is in Appendix A.
@@
$V_{A}(s)$ $=\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}(s)\quad\forall s\in\mathcal{S};$
$V_{B}(s)$ $=\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}(s)\quad\forall s\in\mathcal{S}.$
@@
for all $s\in\mathcal{S}$ in light of Theorem 5.1. Equation 3 shows that the follower’s FPE $\bar{g}=R_{B}(\bar{f},V_{B})$ is the best response to the leader’s FPE $\bar{f}$. However, $R_{B}(f,V_{B})$ is not a best response to an arbitrary $f\in\mathcal{W}_{A}$, except when $V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s)~{}\forall s\in\mathcal{S}$. In addition, determining any meaning from $R_{B}(f,V_{B})$ is difficult. Therefore, $V_{A}^{\bar{f}\bar{g}}$ does not coincides with the SE value function. Moreover, there the performance of the FPE is not guaranteed, which is also empirically demonstrated in Bucarey et al. (2019).
5.2 State-dependent Markov policy and SSE Policy
An attempt to design another operator that has an SE value function as a fixed point is made. From this attempt, we derive a novel sufficient condition for the existence of the SSE policy that is also a necessary condition for a special case. The game class that is guaranteed to have SSE policies by our sufficient condition includes the game class guaranteed by Bucarey et al. (2022).
First, we define a new policy class, state-dependent Markov policy, in which the whole decision rule for each state is determined depending on the state.
Definition 5.2 ((Leader’s) State-dependent Markov policy).
Let $s\in\mathcal{S}$ be a state visited by the agent. A decision rule that the agent follows at $s$ is determined depending on $s$ when the agent follows a state-dependent Markov policy. In other words, let $\hat{\mathbf{f}}:\mathcal{S}\times\mathcal{S}\to\Delta(\mathcal{A})$ be the agent’s state-dependent Markov policy and $\hat{\mathbf{f}}(\cdot;s)\in\mathcal{W}_{A}$ be the agent’s decision rule at $s$. Then, the agent’s action on $s$ is sampled as $a\sim\hat{\mathbf{f}}(s;s)$ and the follower’s best response action is $R_{B}^{*}(s,\hat{\mathbf{f}}(\cdot;s))\in\mathcal{B}$.
Let $\mathcal{W}_{A}^{\mathcal{S}},\mathcal{W}_{B}^{\mathcal{S}}$ be the sets of state-dependent Markov policies for the leader and the follower, respectively. For simplicity, let $\hat{f}_{s}:=\hat{\mathbf{f}}(\cdot;s)$.
State-dependent Markov policies are an extension of the stationary policy concept, such that $\mathcal{W}_{A}^{\mathcal{S}}\supset\mathcal{W}_{A}$. Similar to the stationary policies, we can define the leader’s state value function under the best-response follower as
$V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$ $:=V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s)$
$:=\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\left[\sum_{t=0}^{\infty}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right],$
where $R_{B}^{*}(\hat{\mathbf{f}})\in\mathcal{W}_{B}^{\mathcal{S}}$ is the follower’s state-dependent Markov policy in which the decision rule on the visited state $s\in\mathcal{S}$ is $R_{B}^{*}(\hat{f}_{s})\in\mathcal{W}_{B}$, and $\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}[\cdot]$ is an expectation for the trajectory $\{(s_{t},a_{t},b_{t})\}_{t=0}^{\infty}$, which is generated under $(\hat{\mathbf{f}},R_{B}^{*}(\hat{\mathbf{f}}),\hat{p},\rho)$.
Let the maximization of $V_{A}^{\hat{\mathbf{f}}{\dagger}}$ on $\mathcal{W}_{A}^{\mathcal{S}}$ be defined as
$\hat{V}_{A}^{*}(s):=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$
for each $s\in\mathcal{S}$. Subsequently, Theorem 5.3 shows that $\hat{V}^{*}$ is the unique fixed point of the operator $\hat{T}_{A}$ defined in Equation 5. The proof is in Appendix B.
Theorem 5.3.
Let $\hat{T}_{A}:\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}$ be defined as
$(\hat{T}_{A}v)(s):=\sup_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\quad\forall s\in\mathcal{S}.$ (5)
Then, $\hat{T}_{A}$ is a contraction mapping and its unique fixed point is $\hat{V}_{A}^{*}$.
We define the set of decision rules $R_{A}^{*}(s)$ for each state $s\in\mathcal{S}$ as
$R_{A}^{*}(s):=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (6)
and we suppose $R_{A}^{*}(s)$ is non-empty for all $s\in\mathcal{S}$. Lemma 5.4 provides a way to construct a state-dependent Markov policy by using $R_{A}^{*}$ that realizes $\hat{V}_{A}^{*}$. This proof is in Appendix C.
Lemma 5.4.
Let $\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}$ be a leader’s state-dependent Markov policy that satisfies $\hat{\mathbf{f}}(\cdot;s)\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. Then, $V_{A}^{\hat{\mathbf{f}}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}\hat{V}_{A}^{*}$ holds.
Finally, Theorem 5.5 provides a sufficient condition for the existence of SSE policies. The proof is in Appendix D.
Theorem 5.5.
Let $R_{A}^{*}(s)$ be a subset of $\mathcal{W}_{A}$ defined in Equation 6 for each $s\in\mathcal{S}$. Then, if there exists $\hat{f}\in\mathcal{W}_{A}$ such that $\hat{f}\in\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$, $\hat{f}$ is an SSE policy.
If we assume the sufficient conditions for the existence of SSE policies in Bucarey et al. (2022), $\gamma_{B}=0$ holds, or the transition does not depend on the follower’s action. Then, for $f\in\mathcal{W}_{A}$, because $R_{B}^{*}(s,f)$ depends only on $s\in\mathcal{S}$ and the leader’s action distribution $f(s)\in\Delta(\mathcal{A})$ on state $s$, the part to the right of the $\max$ operator in the right-hand side of Equation 6 depends only on $f(s)\in\Delta(\mathcal{A})$ and not on all $f(s^{\prime})\in\Delta(\mathcal{A})$ such that $s^{\prime}\neq s$. Thus, $R_{A}^{*}(s)$ is a restricted set with respect only to the action distribution $f(s)\in\Delta(\mathcal{A})$, which shows that there exists $f^{\prime}\in\mathcal{W}_{A}$ such that $f^{\prime}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. Then, the existence of SSE policies is guaranteed by Theorem 5.5. This means that the game class whose existence of SSE policies is guaranteed by Theorem 5.5 includes the game class whose existence is guaranteed by Bucarey et al. (2022).
Rather than considering a myopic follower, if we consider the situation that the leader’s discount rate $\gamma_{A}=0$, the sufficient condition in Theorem 5.5 becomes a necessary condition. Its proof is provided in Appendix E.
Corollary 5.6.
If $\gamma_{A}=0$, then an SSE policy exists if and only if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is non-empty.
Even if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is empty, we can construct stationary policies utilizing $R_{A}^{*}$, which is asymptotically obtained using $\hat{T}_{A}$ in light of Theorem 5.3. A possible approach is to choose $f$ contained in $R_{A}^{*}(s)$ for most of the states or to extract one non-dominated by stationary policies in $R_{A}^{*}$. However, the performance is not guaranteed for the stationary policies obtained as aforementioned. This is because any stationary policy not included in $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ does not achieve $\hat{V}_{A}^{*}$, and its performance may be significantly worse than $\hat{V}_{A}^{*}$ as $R_{B}^{*}(f)$; hence, $\hat{V}_{A}^{*}$ is not necessarily continuous with respect to $f$.
6 Policy Iteration for Pareto-Optimal Policies

### v2

@@ 4 Related Work
Bucarey et al. (2022) proposed an algorithm for computing strong SSEs in the 2-player general-sum stochastic games, where strong SSEs are introduced to break ties for the follower’s best response, which corresponds to a specific tie-breaking mechanism for the follower’s best response in our setting. They provide sufficient conditions for their algorithm to converge to the SSEs. However, there is little assurance of the leader’s performance of the obtained policy when it does not converge to the SSEs or when there is no SSE. This problem is because the fixed point of the dynamic programming operator used in Bucarey et al. (2022) is different from the SE value function. We discuss this in detail in Section˜5.
@@
A primary case where the algorithm of Bucarey et al. (2022) converges and the fixed point coincides with the SE values is when the follower is myopic, i.e., when the follower maximizes its immediate rewards. This setting is adopted in some existing SSG studies. For example, Zhao et al. (2023) derived the upper bound of the regret of the leader’s payoff in the cooperative tasks, where the reward function is shared between the leader and the follower, with information asymmetry under the myopic follower. Zhong et al. (2023) proposed an algorithm in the SSG where the group of myopic followers forms a Nash equilibrium among them with analyses on the upper bound of the regret and suboptimality of performance. In contrast to these studies, we focus on the situation where the follower is not myopic.
@@
Zhang et al. (2020a) also considered a similar problem setting, where the reward setting is general-sum and the follower is not myopic. The differences are that the transition function is unknown (i.e., in the RL setting), the follower’s policy takes the leader’s action as input, and only deterministic stationary policies are considered for the leader’s policies. The algorithm proposed by Zhang et al. (2020) shares its principal foundation with that of Bucarey et al. (2022) (see Appendix˜I). This means that they use the same operator, implying that the limitation of the method of Bucarey et al. (2022) applies to that of Zhang et al. (2020).
Policy gradient methods have been proposed for SSGs with non-myopic followers (Zheng et al., 2022; Vu et al., 2022) based on the implicit function theorem (Fiez et al., 2020). They guarantee the convergence to differential SEs (DSEs), which are subsets of local SEs. However, these methods are not applicable in our setting, except when using the surrogate model of the follower, due to their centralized nature of learning. Moreover, the leader’s policy obtained by these algorithms does not necessarily form the DSEs under the best-response follower because the follower’s strategies in DSEs are not always the best response.
Several studies concentrate on the cooperative setting. Kao et al. (2022) and Zhao et al. (2023) analyzed the cooperative tasks with information asymmetry, where the leader cannot observe the follower’s action (Kao et al., 2022) and the reward functions are known only to the follower (Zhao et al., 2023). Kononen (2004) and Zhang et al. (2020a) point out Pareto efficiency and uniqueness as advantages of Stackelberg equilibria to Nash equilibria in cooperative tasks.
We proposed a novel algorithm for general-sum SSGs with the non-myopic best-response follower under the assumptions that the transition function is known and that the follower’s best response is computable in a reasonable time. Unlike the previous methods, iterative methods by the operator (Bucarey et al., 2022; Zhang et al., 2020a) and policy gradient methods (Zheng et al., 2022; Vu et al., 2022), our algorithm guarantees monotone improvement of the leader’s state values under the best-response follower. Furthermore, we give a convergence guarantee even when no SSE policy exists.
5 Analysis on DP Operators
Dynamic programming operators are the core of the solutions of single-agent MDPs and the foundation of RL algorithms (Sutton and Barto, 2018). For stochastic games, the operator is extended as an operation of solving a one-step game by multiple players, and methods for Nash equilibria are proposed with convergence guarantees (Hu and Wellman, 2003).
For SSGs, however, existing methods based on such a one-step game operator (Bucarey et al., 2022; Zhang et al., 2020a) have problems in terms of the leader’s performance in our setting. This section describes it in detail. In summary, we analyze the fixed points of the operator and show that, while the equilibrium formed by the fixed point holds under the best-response follower, the obtained leader’s policy may perform poorly when the fixed point is not the SSE, since monotone improvement in the leader’s performance is not guaranteed.
The one-step game operator $T:\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$ is defined for SSGs as
$(Tv)_{i}(s):=r_{i}(s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))\\ +\gamma_{i}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))}\left[v_{i}(s^{\prime})\right]\quad(i\in\{A,B\}),$ (2)
where $v:=(v_{A},v_{B})\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$,
@@ 5 Analysis on DP Operators
$R_{A}(s,v):=\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},v_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},v_{B}))}\left[v_{A}(s^{\prime})\right],$
@@
which is assumed to exist for all $s$ and vv, [^1] This assumption is necessary to derive a policy corresponding to a value function, and Bucarey et al. (2022) implicitly assumes it. However, it is not guaranteed in general because the inside of $\operatornamewithlimits{argmax}$ is not necessarily continuous with respect to $f_{s}$, even though $\Delta(\mathcal{A})$ is compact. Conversely, $R_{B}$ always exists because $\mathcal{B}$ is finite. and
@@
$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)\\ +\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right],$
@@
which is assumed to be a singleton for all $s$, $f_{s}$, and $v_{B}$. [^2] Bucarey et al. (2022) defines the strong version of $R_{B}$ similarly to the strong SSE. In this paper, we assume that $R_{B}$ is a singleton to simplify the discussion, in accordance with our goal of finding the SSE policies. As the core of our analysis does not depend on the uniqueness of $R_{B}$, the result can be easily extended to the method of Bucarey et al. (2022). Equation˜2 is regarded as a simplified version of the operator defined in Bucarey et al. (2022). The single update of current values $v$ with $T$ as $v_{i}^{\prime}(s)=(Tv)_{i}(s)$ can be seen as solving a normal-form (i.e., one-step) Stackelberg game for player $i$ at given state $s\in\mathcal{S}$ where the payoff functions are given by current Q functions
$q_{i}(s,f_{s},b):=r_{i}(s,f_{s},b)+\gamma_{i}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{i}(s^{\prime})\right]\quad(i\in\{A,B\})$
for the leader’s mixed strategy $f_{s}$ and the follower’s pure strategy $b$. In this context, $R_{B}(s,f_{s},v_{B})$ is the follower’s best response to $f_{s}$, $R_{A}(s,v)$ is the leader’s Stackelberg equilibrium strategy, the pair $(R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))$ is the Stackelberg equilibrium, and $((Tv)_{A}(s),(Tv)_{B}(s))$ is the equilibrium payoffs of the normal-form game for given $s$ and $v$.
@@
Bucarey et al. (2022) proposed a method to construct a stationary equilibrium policy corresponding to the fixed point of the operator, which is called fixed point equilibrium (FPE). If there exists a fixed point $V:=(V_{A},V_{B})$ of $T$, the pair of the value functions of the FPE $(\bar{f},\bar{g})$ defined as, for any $s\in\mathcal{S}$,
$\bar{f}(s)\in R_{A}(s,V),\quad\bar{g}(s)\in R_{B}(s,\bar{f}(s),V_{B})$
coincides with $V$ (Bucarey et al., 2022). Since $T$ is a contraction mapping for $v_{A}$ under certain conditions (Bucarey et al., 2022) [^3] One example is a condition that $r_{A}$ is an affine transformation of $r_{B}$ (i.e., cooperative or competitive with $r_{B}$). In general, $T$ does not always converge, which is empirically demonstrated for the aforementioned example game (Figure 1) in Jean-Marie et al. (2022)., we can obtain the fixed point asymptotically by repeatedly applying $T$ with arbitrary initial values.
@@
Let us consider the characteristics of the fixed point $(V_{A},V_{B})$ of the operator $T$ in general situations where the fixed point exists. Theorem˜5.1 reveals a general property of $(V_{A},V_{B})$. The proof is in Appendix˜A. (Appendices are provided as supplementary material.)
@@
$V_{A}(s)=\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}(s)\kern 5.0pt\forall s\in\mathcal{S};\quad V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}(s)\kern 5.0pt\forall s\in\mathcal{S}.$
@@
for all $s\in\mathcal{S}$ in light of Theorem˜5.1. Equation˜3 shows that the follower’s FPE $\bar{g}=R_{B}(\bar{f},V_{B})$ is the best response to the leader’s FPE $\bar{f}$. However, $R_{B}(f,V_{B})$ is not the best response to arbitrary $f\in\mathcal{W}_{A}$, except when $V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s)~\forall s\in\mathcal{S}$.
The results of Theorem˜5.1 illuminate that the leader’s performance $V_{A}$ of FPEs is the maximum of the leader’s state value on $f\in\mathcal{W}_{A}$ under $R_{B}(f,V_{B})$, but not the maximum under the follower’s best responses $R_{B}^{*}(f)$ like SE value functions. Therefore, the leader’s performance of FPEs at each state can be less than the SE value, which is empirically demonstrated in Bucarey et al. (2019). In addition, it is difficult to determine any meaning from $R_{B}(f,V_{B})$ for arbitrary $f\in\mathcal{W}_{A}$, which makes the performance of FPEs unpredictable.
Although Bucarey et al. (2022) derives the sufficient condition for SSEs to exist and for FPEs to coincide with the SSEs, this condition is not compatible with our problem setting or can be a strong assumption. The sufficient condition is equivalent to the union of two conditions (Bucarey et al., 2022). The first is that the follower’s discount rate $\gamma_{B}=0$, which is a condition on the follower’s learning algorithm, but we admit no other assumptions on the follower than the best responsivity. The second is that the transition function does not depend on the follower’s actions, which can limit the applications of their approach. In the end, finding FPEs does not guarantee the leader’s performance in our setting.
To overcome the problem of SSGs that SSEs do not always exist, in the next section, we propose an algorithm that guarantees monotone improvement in the leader’s performance toward alternative equilibria, Pareto-optimal policies.
6 Pareto-Optimal Policy Iteration